What is cop of carnot engine
1
Expert's answer
2020-12-02T07:31:51-0500

By definition,

"\displaystyle COP = \frac{Q}{W}"

where where Q is the useful heat supplied or removed by the considered system and W is the work required by the considered system.

For Carnot cycle or Carnot engine,

"\displaystyle COP = \frac{T_H - T_C}{T_H} = 1 - \frac{T_C}{T_H}"

"T_C" is the absolute temperature of the cold reservoir, and "T_C" is the absolute temperature of the hot reservoir.